OpenStudy (anonymous):

gr. 11 functions help..

OpenStudy (anonymous):

\[describe how the graphy of y = x ^{2} can be transformed into the graph of this quadratic function... y= 1/4 (x -5)^{2}\]

OpenStudy (anonymous):

what's the rest of that? i didnt catch it

OpenStudy (anonymous):

and what math is this, i only know precal.....

OpenStudy (anonymous):

lemme try typing that again :P ... describe how the graph of y=\[x ^{2}\] can be transformed into the graph of this quadratic function... y= 1/4 \[(x-5) ^{2}\] any ideas?? :S

OpenStudy (anonymous):

let me clarify.. it should be y- 1/4 [(x-5)^{2}\] and how can y= \[x ^{2}\] be transformed (h,k ,a ) to graph the equation above...

OpenStudy (anonymous):

oooh, dear lord. i cannot type today!! sorry the first line should be y= 1/4 \[(x-5)^{2}\]